OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)?

OpenStudy (anonymous):

y + 5 = 1(x + 2) y + 8 = −1(x − 1) y − 5 = −1(x − 2) y − 8 = 1(x − 2)

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

C